### v1

# CAMPOS II. The onset of protostellar disk substructures and planet formation

11institutetext: The NASA Hubble Fellowship Program Sagan Fellow
11email: chenghan.hsieh@utexas.edu
22institutetext: Department of Astronomy, Yale University, New Haven, CT 06511, USA 33institutetext: Department of Astronomy, The University of Texas at Austin, 2515 Speedway, Stop C1400, Austin, Texas 78712-1205, USA 44institutetext: Max Planck Institute for Extraterrestrial Physics, Gießenbachstraße 1, 85748, Garching bei München, Germany 55institutetext: Departamento de Astronomía, Universidad de Chile, Camino El Observatorio 1515, Las Condes, Chile 66institutetext: Department of Physics, Middlebury College, Middlebury, VT 05753, USA 77institutetext: Los Alamos National Laboratory, New Mexico 87545, USA 88institutetext: National Science Foundation-Simons AI Institute for Cosmic Origins, The University of Texas at Austin, 2515 Speedway, Stop C1400, Austin, Texas 78712-1205, USA
CAMPOS II. The onset of protostellar disk substructures and planet formation
Cheng-Han Hsieh {CJK}UTF8bsmi(謝承翰) 112233    Héctor G. Arce 22    María José Maureira 44    Jaime E. Pineda 44    Dominique Segura-Cox 33    Diego Mardones 55    Michael M. Dunham 66    Hui Li{CJK}UTF8gbsn (李晖) 77    Stella S. R. Offner 3388
(Submitted April 15, 2025 )
The 1.3 mm CAMPOS survey has resolved 90 protostellar disks with $\sim$15 au resolution across the Ophiuchus, Corona Australis, and Chamaeleon star-forming regions. To address the fundamental question, ‘When does planet formation begin?’, we combined the CAMPOS sample with literature observations of Class 0-II disks (bolometric temperature, $T_{\rm bol}\leq 1900$ K), all mapped at 1.3 mm with resolutions ranging from 4 to 33 au. To investigate substructure detection rates as a function of bolometric temperature, we restricted the sample to disks observed at the 1.3 mm wavelength, with inclinations below 75, linear resolution $\leq 20$ au and resolved with at least four resolution elements ($\theta_{\rm disk}/\theta_{\rm res}\geq 4$). We also considered the effects of extinction correction and the inclusion of Herschel Space Telescope data on the bolometric temperature measurements, to constrain the lower and upper limits of bolometric temperature for each source. We find that by $T_{\rm bol}\sim 200-400\,K$, substructure detection rates increased sharply to $\sim$60 %, corresponding to an approximate age of $0.2-0.4$ Myr. No substructures are detected in Class 0 disks. The ratio of disk-averaged brightness temperature to predicted dust temperature shows a trend of increasing values toward the youngest Class 0 disks, suggesting higher optical depths in these early stages. Classifying disks with substructures into those with and without large central cavities, we find both populations coexisting across evolutionary stages, suggesting they are not necessarily evolutionarily linked. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage, and the emergence of these structures at $T_{\rm bol}\sim 200-400\,K$ could represent only an upper limit.
Key Words.:
giant planet formation – ALMA Observation – protostellar disks
1 Introduction
One of the major challenges in the field of planet formation is understanding the relationship between circumstellar disk properties and the outcome of the planet formation process (Miotello et al., 2023). The advent of high-resolution observations by the Atacama Large Millimeter/Submillimeter Array (ALMA) has fundamentally transformed our understanding of planet formation, unveiling intricate substructures within Class II disks (age $\gtrsim$1 Myr). These substructures manifest varied natures - some potentially sculpted by pre-existing planets, while others, such as dense rings, may act as nurseries for the formation of planetesimals and subsequent planet generations (ALMA Partnership et al., 2015; Cieza et al., 2017; Pérez et al., 2019; Dong et al., 2017; Jin et al., 2016). The prevalence of substructures in Class II protoplanetary disks has raised questions about how early these substructures form, and thus when the planet formation process begins (Bae et al., 2023).
Previous searches for disk substructures in Class 0/I disks have been limited and focused on the brightest and largest disks (e.g., Sakai et al. 2019; Segura-Cox et al. 2020; Sheehan et al. 2020; Cieza et al. 2021; Ohashi et al. 2023; Lee et al. 2023). Our recent ALMA CAMPOS program has uniformly surveyed 18 Class 0, 31 Class I, 22 flat-spectrum, and 19 early Class II protostellar disks with a resolution of 15 au in nearby star-forming regions within 200 pc (Hsieh et al., 2024). Together with existing literature data, forming a total sample of 116 disks, we conducted a statistical study of the onset of disk substructures in embedded protostellar disks with a large sample and in different environments.
This paper is organized as follows: In Section  2 and  3, we describe the substructure identification, mass calculation using the disk fluxes, and the construction of a homogeneous sample to search for the onset of disk substructures. We restricted the disk sample observed at the same 1.3 mm wavelength, with inclination angle below 75 and resolved at least 4 beams. In Section  4, we present the main result of this paper: the substructure detection rate statistics as a function of evolutionary stage tracked by bolometric temperature. In addition, we discuss the existence of 2 populations of disk substructures. We discuss our results in Section  5, and present our conclusion in Section  6.
2 Data
2.1 Observational data and the sample construction
The detailed data reduction and the survey design for the CAMPOS survey are described in Hsieh et al. (2024). We combined the CAMPOS uniform survey of 90 protostellar disks at $\sim 14-18$ au in Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus (Hsieh et al., 2024) with protostellar disks from the high-resolution ($\sim 7$ au) eDisk survey (Ohashi et al., 2023), medium-resolution ($\sim 28$ au) ODISEA survey (Cieza et al., 2019), high-resolution ($\sim 4$ au) ODISEA survey (Cieza et al., 2021), and high-resolution ($\sim 4$ au) DSHARP survey (Andrews et al., 2018a) to study the onset of disk substructures. We also included the high resolution ($\sim 4$ au) observation of HL Tau (ALMA Partnership et al., 2015).
[…]
Related to the resolution, we note that the absolute linear resolution naturally determines the smallest substructure size/scales that can be probed. The smallest detectable substructure scale is around half the beam size (Martí-Vidal et al., 2012). We would use this to discuss the statistics of the size of substructures in the sample.
Besides the effective angular resolution and linear resolution, it is also important to compare observations at the same wavelength. The thermal and scattered continuum emission at a wavelength $\lambda$ is dominated by grains with a size of $\sim\lambda/2\pi$ (Mie, 1908), meaning ALMA 1.3 mm observations, such as the CAMPOS survey as well as DSHARP and eDISK, are more efficient at probing few 100 $\mu$m sized grains. In contrast, disk observations in the optical/near-infrared mainly probe much smaller sub-micrometer particles. To ensure a consistent survey probing the substructures from similar solids, we only include protostellar disks observed at 1.3 mm wavelength by ALMA. Finally, the occurrence rate of substructures is also influenced by the disk’s inclination angle. Due to the increasing optical depth and reduced emitting size, searching for disk substructures in edge-on disks is extremely challenging. While most of the high-resolution observations towards the more evolved Class II disks avoided highly inclined sources (e.g., DSHARP), the population of Class 0/I disks in nearby regions is more limited, and thus samples of very high-resolution observations (e.g., eDisk) contain a significant amount of highly inclined disk. To take this into account, we exclude all disks in the sample with inclination angles larger than 75.222We assume thin circular disks and use the major-to-minor axis ratio of the dust continuum data to compute the inclination angle ($i$), that is $i=sin^{-1}(R_{\rm min}/R_{\rm maj})\times 180^{\circ}/\pi$.
2.2 Substructure criteria
In this paper, we labeled disks with substructures if the disk has rings, gaps, a central cavity, or spirals. Sources with asymmetric intensity profiles but lacking annular or spiral structures (e.g., IRAS 16293-2422B, Zamponi et al. 2021, 2024; Hsieh et al. 2024) are counted as disks without substructures.
We note that following the above criteria, all of the disks considered as with substructures in this work have been already confirmed as such in separate published papers (CAMPOS: Hsieh et al. (2024), eDisk: Ohashi et al. (2023), ODESIA: Cieza et al. (2019, 2021), DSHARP: Andrews et al. (2018a).), and all correspond to substructures that are immediately visible in the 1.3 mm image. Sources that show small variations as a function of distance in $uv-$space or small variation in the intensity radial profile but appear smooth in the image plane counted here as disk without substructures (e.g., Michel et al. 2023). This is a conservative approach, and we may miss some sources with borderline detections, introducing a bias against younger disks with smaller and lower-contrast (shallower) substructures (e.g., Maureira et al. 2024a). The implications of this will be further discussed in Section 5.
3 Analysis
3.1 Disk fluxes and sizes
[…]
3.2 Deriving the mass of protostellar disks
We adopted the method outlined in the ODISEA survey of Ophiuchus protoplanetary disks to convert the 1.3 mm disk flux into dust mass. This method assumes optically thin emission, a constant temperature of 20 K and an opacity coefficient of $\kappa_{\nu}=(\nu/100\,{\rm GHz})$ cm2 g-1 (Williams et al., 2019). We then assumed a gas-to-dust ratio of 100 to obtain the disk mass (Bae et al., 2023). We caution that disks might not be fully optically thin, particularly at the early stages (e.g., Li et al., 2017; Zamponi et al., 2021; Lin et al., 2021; Maureira et al., 2022), but also at the Class II stage (Andrews et al., 2018b; Tazzari et al., 2021; Chung et al., 2024). In this case, the estimated masses can be conservatively considered a lower limit. The protostellar disk mass measurement is for better comparison with other works and as a guide, and it does not affect the analysis of the onset of disk substructures.
3.3 Bolometric temperature as a disk evolutionary tracer
We cross-matched all protostellar disks with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al., 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog (Riwaj Pokhrel, private communication), to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems. The eHOPS catalog contains 1-850 $\mu$m Spectral Energy Distributions (SEDs) assembled from Two Micron All Sky Survey (2MASS), Spitzer, Herschel Space Telescope (Herschel), Wide-field Infrared Survey Explorer (WISE), and James Clerk Maxwell Telescope (JCMT)/SCUBA-2 data. The first eHOPS paper for the Serpens and Aquila molecular clouds was published by Pokhrel et al. (2023). For all other clouds, the SED and protostellar system properties are available in the NASA/IPAC Infrared Science Archive. For sources without a counterpart in both eHOPS and Dunham et al. (2015), we adopted the $T_{\rm bol}$ from the literature. Given that we want to investigate the emergence of substructures, we only included samples with $T_{\rm bol}\leq 1900$ K. The final list of sources with $T_{\rm bol}\leq 1900$ K and an inclination angle less than $75^{\circ}$ is shown in Table LABEL:table:substructure_data. In Table LABEL:table:CAMPOS_Tbol, we present the alternative names for each source in the CAMPOS survey for the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources. The cross-matched results for the ODISEA survey are presented in Table LABEL:table:ODISEA2019_Tbol.
In this paper, we consider the bolometric temperature measurements from both the eHOPS catalog and the Dunham et al. (2015) catalog to trace the onset of protostellar disk evolution. Compared to Dunham et al. (2015), the eHOPS catalog provides more complete spectral energy distributions (SEDs) from $1$ to $850\,\mu$m. However, the infrared portion of the SED is not corrected for extinction (Pokhrel et al., 2023). To maintain consistency with the eHOPS catalog, we use the bolometric temperatures without extinction correction from Dunham et al. (2015) for sources without a counterpart in eHOPS. Extinction correction increases the infrared flux and increases the bolometric temperature. Therefore, results based on the eHOPS bolometric temperatures represent a lower age limit for the onset of disk substructures. On the other hand, to investigate the upper age limit for the onset of disk substructures, we adopt the extinction-corrected bolometric temperature from Dunham et al. (2015). The bolometric temperatures were derived from the SED fitting without Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). The extinction correction applied to the infrared data increases the infrared flux and thus further increases the bolometric temperature (Dunham et al., 2015). Results based on the extinction-corrected bolometric temperature from Dunham et al. (2015) represent the upper age limit for the onset of disk substructures as longer-wavelength Herschel data are not included from the SED fitting to compute $T_{\rm bol}$.
The bolometric temperature measured from the eHOPS catalog and Dunham et al. (2015) differs from source to source. The decrease in $T_{\rm bol}$ is not uniformly large when including Hershel data into the Dunham et al. (2015). It depends both on the shape of the SED and on how well the SED was sampled by Dunham et al. (2015). While Herschel photometry was not available for the SED fits in Dunham et al. (2015), many of the protostars had 450 $\mu$m observations from SCUBA, some had 350 $\mu$m observations from SHARC-II on Caltech Submillimeter Observatory, and some brighter objects had Infrared Astronomical Satellite (IRAS) or Infrared Space Observatory (ISO) far-infrared photometry. Spitzer 70 $\mu$m photometry was also available for nearly the full Spitzer-GB sample, but at a factor of 3 lower spatial resolution than Herschel, resulting in confusion for some closer protostars. As a result, some $T_{\rm bol}$ values change by a large amount, while some change by a very small amount. We present the $T_{\rm bol}$ from both eHOPS and Dunham et al. (2015) to quantify the range for the onset of protostellar disk substructures.
4 Results
4.1 Protostellar disk substructures versus $T_{\rm bol}$ and disk properties
Figure 1: Protostellar disk substructure detections as a function of disk radius, mass, and bolometric temperature. The latter is taken from the eHOPS catalog. The gray dashed lines indicate the boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). While the $T_{\rm bol}$ is not linear with age, in general the Class 0 sources have $T_{\rm bol}<70$ K, Class I sources have $65070<T_{\rm bol}<$ 650 K, and Class II and Class III sources have $T_{\rm bol}>650$ K (Bouvier et al., 2021). The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
Figure 2: Substructure detection as a function of effective resolution, 1.3 mm flux and bolometric temperature, the latter from the the eHOPS catalog. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
Figure 3: The upper limit of the onset of disk substructures. Same as Figure 1, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infared data, which further increases the bolometric temperature (Dunham et al., 2015). The effective resolution and the ALMA Band 6 1.3 mm flux of protostellar disks are color-coded and shown in different symbols according to the presence or absence of substructure. Compact disks with diameters smaller than 4 beams are indicated by crosses. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
Figure 4: The upper limit of the onset of disk substructures. Same as Figure 2, but the bolometric temperature is derived from the SED fitting without the Herschel Space Observatory data ($70-500\,\mu$m) (Dunham et al., 2015). Extinction correction is applied to the infared data, which further increases the bolometric temperature (Dunham et al., 2015). The crosses represent compact disks that are not well-resolved with diameters less than 4 times the angular resolution of the observation (beams). The lower limit of the onset of the disk substructures at 200 K found in Figure 1 is also labeled. The black arrow is used to indicate that the particular source has also been classified as Class II in the literature (Narayanan et al., 2023).
Figure 1 illustrates the radius and mass of the protostellar disks as a function of bolometric temperature ($T_{\rm bol}$) derived from the eHOPS catalog. The black dashed lines indicate the $T_{\rm bol}$ that marks the boundary between the Class 0 and Class I, and between the Class I and Class II evolutionary phases (Chen et al., 1995), along with the corresponding expected duration of each Class (Dunham et al., 2015). For reference, we also include the sources that do not comply with the effective resolution criteria ($\theta_{D}/\theta_{\rm res}\geq 4$). The latter sources have, in general, smaller sizes and fluxes/masses.
The earliest substructure observed, IRAS 04169+2702, corresponds to a disk with an approximate age of $0.1-0.2$ Myr ($T_{\rm bol}\sim 163$ K). If we consider the lower limit of $T_{\rm bol}$ of CFHTWIR-Oph 79 ($T_{\rm bol}\sim 110$ K), then the earliest substructure could be even younger. [^4] The lower limit of the $T_{\rm bol}$ from this source corresponds to the calculation from eHOPS. On the other hand, in the literature, this source has been classified as a Class II, hence the lower limit.
We find no detection of substructure for all the Class 0 disks in the sample. Figure 2 shows the effective resolution and flux of the disks with and without substructures as a function of bolometric temperature, color-coded by the linear resolution. We can observe that the lack of substructure in Class 0 disks is true across resolutions from 4 to 18 au.
Figure 3 and 4 show similar plots as Figure 1 and 2 but considering the upper limit to bolometric temperatures from Dunham et al. (2015). In that case, the earliest substructures discussed above correspond to $T_{\rm bol}$ of 163 and 300 K, for IRAS 04169+2702555For IRAS 04169+2702, the $T_{\rm bol}$ is adopted from Ohashi et al. (2023). and CFHTWIR-Oph79, respectively.
4.2 Disk substructure detection rates
Figure 5: Disk substructure detection rates versus the evolutionary stage traced by bolometric temperature. The blue and orange represent the lower and upper limits with the bolometric temperature ($T_{\rm bol}$) measured by the eHOPS (with Herschel, but not extinction corrected) and by Dunham et al. (2015) (no Herschel, but extinction corrected), respectively. The bin size of 8 disks is used to compute the substructure detection rates. The shaded region represents the $1\sigma$ uncertainty from counting statistics. All protostellar disks are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au.
We present the protostellar disk substructure detection rates as a function of bolometric temperature in Figure 5. Each disk is assigned a value of 1 if it exhibits substructures and 0 if it does not. We compute a moving average with a window size of 8, where the kernel slides through $T_{\rm bol}$ and calculates the mean for every group of 8 disks. This mean represents the substructure detection rate. Due to the fixed window size, the detection rates are sampled at discrete levels of 0/8, 1/8, 2/8, …, up to 8/8 — effectively producing 8 bins in the substructure detection rate space.
All protostellar disks in Figure 5 are observed by ALMA Band 6 at 1.3 mm wavelength, have a disk diameter greater than 4 times the resolution ($\theta_{D}/\theta_{res}\geq 4$), with an inclination angle less than 75 , and linear resolution less than 20 au. We adopt the moving average with a window size of 8 to ensure enough sampling resolution across $T_{\rm bol}$ and substructure detection rate space. The blue and orange represent the case when the bolometric temperature is adopted from eHOPS catalog or Dunham et al. (2015). The shaded region represents the uncertainty from the binomial counting uncertainty ($\sigma_{b}$), which is calculated as
$\sigma_{b}=\sqrt{\frac{r(1-r)}{n}},$ (1)
where $r$ is the substructure detection rate, and $n=8$ is the total number of sources in the window. The substructure detection rate adopting the Dunham et al. (2015) $T_{\rm bol}$ in general agrees with the results adopting the eHOPS $T_{\rm bol}$, with a delay of $\sim 200$ K in the rapid rise of disk substructures detection rate at $T_{\rm bol}\sim 200-400$ K.
Figure 5 shows that the rate quickly reaches $\sim$60at $T_{\rm bol}$ of 200–400 K, clearly indicating that substructure during the Class I stage is not rare. The Class I protostellar disk substructure detection rate in Figure 5 is also consistent with the $\sim$50substructure detection rate in a sample of 62 Class II protoplanetary disks observed with an effective resolution $3<\theta_{D}/\theta_{res}\leq 10$ (Bae et al., 2023), strongly suggesting that the current detection rate of 60% from $T_{\rm bol}$ 200–400 K, is only a lower limit given the more limited resolution of the younger disk sample (see Figure 2). The high Class I disk substructure detection rate supports that the onset of disk substructures does not occur when disks transit into the Class II phases, but instead firmly into the Class I stage and possibly earlier (Section 5).
Figure 6: From our CAMPOS survey, we divide the disk substructures based on the presence or absence of a central cavity (Hsieh et al., 2024). The “Cavity-type” disk substructure, or “C-type” shows a large central cavity. The second type of disk substructures is the “Centrally-filled type” disk substructure, or “F-type”. Both types of disk substructures can be detected throughout the Class I to Class II phases.
Figure 7: Most of the multiple-ring systems are centrally filled (Type F) (Andrews et al., 2018a; Cieza et al., 2021; Ohashi et al., 2023; Hsieh et al., 2024; Cieza et al., 2019; ALMA Partnership et al., 2015; Stephens et al., 2023). The lack of a large central cavity ($\geq$ 25 % of the disk diameter) implies that a centrally filled disk is a general feature of multiple-ring systems. The cyan circle represents the beam size. The white lines mark the scale of 50 au. Note that HD 163296, RU Lup, AS 209, GW Lup, HD 142666, HD 143006, and Sz 129 either do not have reliable bolometric temperature ($T_{\rm bol}$) measurements or their $T_{\rm bol}\geq 1900$ K, thus are not included in this study.
4.3 The protostellar disk substructures with and without central cavities
We found that some protostellar disks have a central cavity in Class I and flat-spectrum phase. To further investigate this, we divide the substructure in two more general groups. The cavity-type (C-type) and the centrally filled type (F-type). The cavity-type disk substructures exhibit large ($>$10 au) cleared central dust cavities. These C-type protostellar disk substructures display morphologies similar to the more evolved Class II/Class III transition disks (van der Marel et al., 2016b). The origin of the central dust cavity in transition disks can be attributed to a range of processes, like extensive particle growth (Tanaka et al., 2005; Dullemond & Dominik, 2005), photoevaporative mass-loss in winds (Clarke et al., 2001), dead zones (Regály et al., 2012) or embedded stellar or planetary companions (Ireland & Kraus, 2008). In particular, in systems with super-Jovian planets or stellar binaries that can clear the material in disks, deep gas gaps are expected inside cavities in the disk dust distribution (Price et al., 2018; Muley et al., 2019; van der Marel et al., 2021). The other type of disk substructure population is the centrally-filled type (F-type) disk substructure. In contrast to the C-type disk substructures, F-type disk substructures do not exhibit a central cavity.
Figure 6 shows that these 2 types of substructures as a function of evolutionary state. C-type disk substructures are found across all evolutionary stages from Class I to Class II. The detection of large cavities in younger Class I and Flat-spectrum sources cannot be explained by photoevaporative mass-loss in winds (Clarke et al., 2001). Instead, the central cavity in Class I can possibly be explained by radial drift (Birnstiel et al., 2010; Cridland et al., 2017). The dust particles will experience a headwind due to the gas drag, drifting towards the central protostar (Cridland et al., 2017). Another explanation for these central cavity is tidal interactions with companions. There is no evidence of binarity for any of the disks in Figure 6. We can rule out all binaries with orbital separations greater than 15 au, which corresponds to the observational linear resolution. Protostellar disks with a C-type disk substructures are good candidates to search for gap-opening protoplanets or close protobinaries.
Notably, we found that multiple-ring systems, which have at least two rings with similar brightness within a factor of two at different disk radii, are almost always centrally filled (F-type). Among the 13 multiple-ring systems complied from the literature (Andrews et al., 2018a; Cieza et al., 2021; Segura-Cox et al., 2020; Hsieh et al., 2024; ALMA Partnership et al., 2015; Stephens et al., 2023; Huang et al., 2018), we found that the majority (10 out of 13 or $77\,\%$) are centrally filled (F-type), as shown in Figure 7. The three exceptions (HD 142666, HD 143006, and Sz 129) are among the oldest systems (Huang et al., 2018). Notably, the central cavities in these systems are significantly smaller than their disk sizes (Clarke et al., 2001). Multiple-ring systems with a large central cavity ($\geq 25$ of the disk diameter) are absent or very rare. The high fraction of F-type multiple-ring systems implies that a centrally filled disk is a general feature of multiple-ring systems.
5 Discussion
Figure 8: Left: The brightness temperature of the protostellar disks calculated from the full Planck equation as a function of the evolutionary stage (bolometric temperature). Right: The transition between disks with substructures and without substructures at $T_{\rm bol}\sim 200$ K is not due to the optical depth. We do not see a discontinuity for the brightness temperature to dust temperature ratio ($T_{\rm bright,Planck}\,/\,T_{\rm dust}$) at 200 K. For optically thick disks the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio should be close to 1. We adopted the dust temperature from Equation 1, and excluded the circumbinary disks.
Since the discovery of rings and gaps in HL Tau protoplanetary disk (ALMA Partnership et al., 2015), many more evolved protoplanetary disks substructures (rings, gaps, spirals) have been detected (e.g., Long et al. 2018; Tsukagoshi et al. 2019; Pinilla et al. 2017; van der Plas et al. 2017; van der Marel et al. 2016a; Kudo et al. 2018; Pinilla et al. 2018; Guzmán et al. 2018; van Terwisga et al. 2018; Fedele et al. 2017; Leiendecker et al. 2022; Booth et al. 2024; Shi et al. 2024; Huang et al. 2024; Guerra-Alvarado et al. 2025). Guerra-Alvarado et al. (2025) recently observed over 73 Class II disks in the Lupus molecular cloud at a resolution of 4 au, and found that the substructure detection rate strongly depends on the effective resolution. For disks with sizes between 8 and 30 au ($2\leq\theta_{\rm D}/\theta_{\rm res}\leq 7.5$), the substructure detection rate is approximately 40 %, whereas it increases to $\sim$80for disks larger than 30 au ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) (Guerra-Alvarado et al., 2025). Although many Class II disks are compact (Guerra-Alvarado et al., 2025), the majority of Class II protoplanetary disks observed at high linear (a few au) and effective resolution ($\theta_{\rm D}/\theta_{\rm res}\geq 7.5$) exhibit substructures (Andrews et al., 2018a; Huang et al., 2018; Bae et al., 2023; Guerra-Alvarado et al., 2025). The high substructure detection rate in well-resolved Class II disks indicates that substructures are already widespread by the Class II stage (Bae et al., 2023).
The quest to search for disk substructures in the Class 0/I protostellar phase has generated a great interest in the disk community (Aso et al., 2021; Xu et al., 2023; Lee et al., 2023; Maureira et al., 2024b; Michel et al., 2023; Ohashi et al., 2023). In particular, the recent ALMA Large Program, eDisk, has made a significant contribution, surveying around 20 embedded Class 0/I protostellar disks and found three of the disks in their sample have substructures (L1489 IRS, Oph IRS 63 and IRAS 04169+2702) (Ohashi et al., 2023). Among these, substructures in Oph IRS 63 and L1489 IRS had been previously reported (Segura-Cox et al., 2020; Ohashi et al., 2022). Based on the prevalence of Class 0/I disks without substructures (detection rate 3 out of 20), Ohashi et al. (2023) concluded that substructures form rapidly near the Class I/II transition.
Our analysis of 90 protostellar disks from CAMPOS, incorporating observations from eDISK, ODISEA, and DSHARP, shows that detection rates for Class I and Class II disks become comparable at $T_{\rm bol}\sim 200-400$ K. In this study, we accounted for potential biases, including variations in observational resolution and source inclination. Lower resolution and higher inclination can artificially reduce the visibility of disk substructures, disproportionately affecting the detection of younger disks in current high-resolution surveys. We also consider the possible systematic biases, such as the lack of extinction correction or Herschel data, on the bolometric temperature of each source. The onset of disk substructure at $T_{\rm bol}\sim 200-400$ K, corresponds to an approximate age of $\sim$0.2-0.4 Myr, which is much earlier than the age of $\sim 1\,$Myr – the expected age for when the transition between the Class I and Class II takes place (Evans et al., 2009; Dunham et al., 2015).
It is important to note that the onset of disk substructures in this study could be an upper limit, and the actual onset of disk substructures may occur much earlier, within the Class 0 phase. We took a conservative approach and did not count sources that show small power variations as a function of distance in $uv-$space or small variations in the intensity radial profile but appear smooth in the image plane as disks with substructures (e.g., TMC-1A spiral [^6] We did not include TMC-1A as a substructure source because it was not identified as a disk with a spiral substructure in the eDisk survey (Ohashi et al., 2023), in contrast to Xu et al. (2023).: Aso et al. 2021; Xu et al. 2023, HH211 spiral [^7] HH211 is not included in Table 1, because it is an edge-on source.: Lee et al. (2023), Oph A SM1 gap [^8] Oph A SM1 is not included in Table 1 because it does not have a bolometric temperature measurement.: Maureira et al. (2024b), faint rings in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3999Ring-like structures are not visually identifiable by the naked eye in Oph-emb-6, Oph-emb-9, and GSS 30 IRS 3 (Michel et al., 2023; Hsieh et al., 2024).: Michel et al. (2023)). Our conservative approach minimizes the chances of false positives, but we may miss lower-contrast (shallower) substructures in younger disks with borderline detection. Furthermore, another possible reason for this onset being an upper limit is the choice to include all disks with $\theta_{D}/\theta_{\rm res}\geq 4$, since the detection rate is only $\sim 50\%$ for values between $3<\theta_{D}/\theta_{\rm res}\leq 10$ (Bae et al., 2023).
In addition, optical depth can also affect the detection rate for the younger Class 0 or early Class I disks (e.g., Maureira et al., 2024a). To investigate this possibility further, we estimated the optical depth at 1.3 mm of our disks by computing the ratio between brightness temperature ($T_{\rm B,Planck}$) and the dust temperature. We estimated the brightness temperature of the disk from the average disk intensity using the full Planck equation. The estimated brightness temperature as a function of evolution stage ($T_{\rm bol}$) is plotted in the left panel of Figure 8. We then compared it with the dust temperature given by:
$T_{\mathrm{dust}}=43\left(\frac{L_{\mathrm{bol}}}{1L_{\odot}}\right)^{0.25}$ (2)
, where $L_{\mathrm{bol}}$ is the bolometric luminosity (Tobin et al., 2020). The optical depth ($\tau_{\nu}$) follows the following relationship:
$T_{\mathrm{bright,Planck}}=T_{\mathrm{dust}}\left(1-e^{-\tau_{\nu}}\right)$ (3)
For optically thick disks, the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio is expected to be close to 1. In Figure 8 right panel, we plot the $T_{\rm bright,Planck}\,/\,T_{\rm dust}$ ratio as a function of evolutionary stage.
Although we do not find a sharp transition in optical depth between Class 0 and I, there is an indication of optical depth increasing from Class I back into the Class 0 stage. This supports that the zero detection rate in our large sample for $T_{\rm bol}\leq 163$ K might be affected by high optical depths at 1.3 mm. Further suggestive of this is the fact that the disks with detected substructures in our work seem to have, on average, less estimated optical depth than all the remaining disks with no substructures during the Class 0 and I stages. If the Class 0 protostellar disks are optically thick, then the onset of protostellar disk substructures should occur even earlier.
The rapid formation of disk substructures at early Class I phases suggests that protostars and protoplanets grow and evolve together. During the Class 0 and early Class I phases, the central protostar and protostellar disks are constantly fed by anisotropic infall via streamers (Pineda et al., 2020; Valdivia-Mena et al., 2022; Pineda et al., 2023). This anisotropic infall can perturb the protostellar disks, generating vortices and pressure bumps, resulting in substructure formation (Kuznetsova et al., 2022). It is possible that these earliest disk substructures are triggered by anisotropic infall (Kuznetsova et al., 2022; Zhao et al., 2025). Other mechanisms for substructure formation have also been proposed (see Bae et al. (2023) for a review). While we do not know the origin of these early disk substructures, these structures can possibly trigger subsequent giant planet formation (Lau et al., 2024). The early onset of protostellar disk substructures at $\sim$0.2 – 0.4 Myr is also consistent with recent population synthesis models, which show that rapid substructure formation within 0.4 Myr is needed to explain the observed spectral index distribution in Lupus protoplanetary disks (Delussu et al., 2024). Our findings of the early formation of disk substructures at early Class I phases confirm that star formation sets the initial conditions for planet formation in protostellar disks.
6 Conclusion
In summary, our ALMA study reveals that by $T_{\rm bol}\sim 200-400\,K$ protostellar disk substructure detection rates increased sharply to $\sim$60 %. Our statistical analysis confirms that substructures similar to those in Class II disks are already common by the Class I stage. We argue that this emergence is likely only an upper age limit, as our study suggests increasing optical depth toward the earliest stages.
The prevalence of substructures in the early protostellar stage suggests that planetesimals and giant planets form much more efficiently and earlier than predicted by traditional models (e.g., Pollack et al. (1996)). This early planet formation scenario is also supported by a low abundance of water in the inner Solar System (Morbidelli et al., 2016; Hyodo et al., 2019), the existence of several isotopic reservoirs in the Solar system (Kruijer et al., 2017) as well as the distribution of the calcium-aluminum-rich inclusions (Desch et al., 2018; Brasser & Mojzsis, 2020). Additionally, older protoplanetary disks (around Class II sources) might not have enough dust mass to form planetary systems (Manara et al., 2018).
Our study has pushed the onset of planet formation to the much earlier protostellar phase, in $\sim$0.2–0.4 Myr disks. During this time period, the central protostar still has a high accretion rate and is building up its mass.
Data Availability
All data used in this study are public and can be downloaded from the ALMA Science Archive.
Software
Astropy: Greenfield et al. (2013); Astropy Collaboration et al. (2018, 2022), CASA: McMullin et al. (2007) SciPy: Virtanen et al. (2020)
Acknowledgements.
C.H.H. and H.A are supported in part by NSF grants AST-1714710. C.H.H. is supported by NASA Hubble Fellowship Program under award HF2-51556. D.S.C. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2102405. Special thanks to e-HOPS group, Dr. Riwaj Pokhrel, and Dr. Thomas Megeath for providing the unpublished e-HOPs catalog for protostar classification. C.H.H. thanks Dr. Sean Andrews for the suggestion of considering the effective resolution in constructing a homogeneous sample for this statistical study. This paper makes use of the following ALMA data: ADS/JAO. ALMA #2016.1.00484.L, #2018.1.00028.S, #2019.1.00261.L, #2019.1.01792.S. ALMA is a partnership of the ESO (representing its member states),NSF(USA),and NINS (Japan), together with the NRC (Canada), NSC and ASIAA(Taiwan), and KASI(Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by the ESO, AUI/NRAO, and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.
Appendix A Supplementary Materials
In Table LABEL:table:substructure_data, we present all the disks in this study complied from the CAMPOS survey (Hsieh et al. 2024), eDisk survey (Ohashi et al. 2023), DSHARP survey (Andrews et al. 2018a), and ODISEA survey (Cieza et al. 2021, 2019). All disks have a bolometric temperature less than 1900 K, and inclination angles less than $75^{\circ}$. The Table consists of a homogeneous survey of 5 molecular clouds, Corona Australis, Chamaeleon I & II, Ophiuchus & Ophiuchus North, with additional sources in Taurus, Lupus I and BHR 71 from the literature.
In Table LABEL:table:CAMPOS_Tbol, we cross-matched all the Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources in our CAMPOS survey with the Young Stellar Objects catalog from the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al. 2015), as well as the Extension of HOPS Out to 500 ParSecs (eHOPS) catalog to obtain the bolometric temperature ($T_{\rm bol}$), which serves as a proxy for relative evolutionary age for embedded protostellar systems. The eHOPS catalog contains 1-850 $\mu$m SEDs assembled from 2MASS, Spitzer, Herschel, WISE, and JCMT/SCUBA-2 data. It represents the latest and most reliable SED fitting to date. The first eHOPS paper, covering the Serpens and Aquila molecular clouds, was published by Pokhrel et al. (2023). For all other clouds, the SED and the protostellar system properties are available in NASA/IPAC Infrared Science Archive.
In Table LABEL:table:ODISEA2019_Tbol, we cross-matched all the sources from the ODISEA survey (Cieza et al. 2019) with the Spitzer Space Telescope “cores to disks” (c2d) and “Gould Belt” (GB) Legacy surveys (Dunham et al. 2015), as well as the eHOPS survey (Pokhrel et al. 2023), to obtain bolometric temperature and luminosity. Nine protostellar disks have bolometric temperature measurements but are unresolved in the ODISEA survey (Cieza et al. 2019) and are therefore not included in Table LABEL:table:substructure_data.
Table 1: Protostellar Disk Substructures Table
Name Class eHOPS Dun+15 ${\rm L}_{\rm bol}$ Mass $R_{90\%}$ dist $F_{1.3mm}$ Type Beam Res $T_{\rm br,P}$ $T_{\rm dust}$ Ref
${\rm T}_{\rm bol}$ [K] ${\rm T}_{\rm bol}$ [K] [$L_{\odot}$] [$10^{-5}M_{\odot}$] [au] [pc] [mJy] [mas] [au] [K] [K]
Ophiuchus & Ophiuchus North
VLA 1623B 0 30 30 2.87 21.48 37.3 144 114.2 NA 130 x 100 16.4 19.7 56.0 1
VLA 1623Aa 0 30 30 2.87 10.7 18.5 144 57.1 NA 130 x 100 16.4 35.2 56.0 1
VLA 1623Ab 0 30 30 2.87 9.09 15.9 144 48.3 NA 130 x 100 16.4 39.5 56.0 1
VLA1623A 0 30 30 0.45 39.22 209.7 144 208.5 C 130 x 100 16.4 4.1 35.2 1, 2, 3
circumbinary
IRAS 16293-2422A 0 31 45 16.25 152.08 66.9 144 808.6 NA 62 x 50 8.0 37.7 86.3 1, 4
IRAS 16293-2422B 0 31 45 16.25 229.02 41.3 144 1217.8 NA 62 x 50 8.0 133.9 86.3 1, 4
CB 68 SMM 1 0 38 39 0.89 10.75 26.5 151 52.0 NA 36 x 27 4.7 15.3 41.8 1, 5
Oph-emb 1 0 42 36 0.19 2.26 13.1 144 12.0 NA 73 x 48 8.2 17.5 28.4 1, 5
eHOPS-oph-20b 0a 57 830 1.77 0.54 12.1 144 2.9 NA 130 x 100 16.4 7.8 49.6 1
eHOPS-oph-20c 0a 57 830 1.77 0.48 10.8 144 2.5 NA 130 x 100 16.4 8.2 49.6 1
Oph-emb 6 0 67 120 0.35 8.18 44.0 144 43.5 NA 130 x 100 16.4 8.3 33.1 1
Oph-emb 16 I 74 420 30.02 2.82 5.7 144 15.0 NA 130 x 100 16.4 87.8 100.7 1
Oph-emb 9 I 76 200 0.33 8.05 30.2 144 42.8 NA 130 x 100 16.4 13.1 32.6 1
Oph-emb 10a I 78 180 10.25 9.12 11.0 144 48.5 NA 130 x 100 16.4 77.9 76.9 1
Oph-emb 10b I 78 180 10.25 1.52 6.6 144 8.1 NA 130 x 100 16.4 38.6 76.9 1
ISO-Oph 31 I 97 250 1.7 21.42 64.5 138 124.0 NA 68 x 45 7.6 7.8 49.1 5
Oph-emb 8 I 97 250 21.16 2.27 6.9 144 12.1 NA 130 x 100 16.4 50.9 92.2 1
CFHTWIR-Oph 79 I 112 300 7.78 2.67 46.9 144 14.2 C 56 x 39 6.7 4.6 71.8 1, 6
Oph-emb 12 I 115 230 0.67 0.72 6.3 144 3.9 NA 130 x 100 16.4 22.5 38.9 1
Oph-emb 13 I 129 280 11.11 1.98 4.5 144 10.5 NA 130 x 100 16.4 99.9 78.5 1
Oph-emb 4 I 135 160 0.1 2.43 32.0 144 12.9 NA 130 x 100 16.4 6.2 24.2 1
Oph-emb 11 Flat 163 260 0.71 1.97 16.1 144 10.5 NA 130 x 100 16.4 11.9 39.5 1
ODISEA C4 064 Flat 185 1100 0.026 6.67 39.2 144 35.5 NA 280 x 190 33.2 8.5 17.3 7
IR Cha INa4 I 186 430 0.06 4.43 31.4 179 15.2 NA 120 x 90 18.6 8.6 21.3 1
ODISEA C4 102 I 207 207 0.13 7.43 37.6 144 39.5 NA 280 x 190 33.2 9.4 25.8 7
IRAS 16237-2428 Flat 221 380 1.88 12.7 15.2 144 67.6 NA 130 x 100 16.4 58.3 50.4 1
ISO-Oph-54 I 224 470 0.14 11.0 119.0 141 61.0 F 58 x 38 6.6 3.9 26.3 1, 8
IRAS 16442-0930 I 230 280 0.44 19.96 37.5 144 106.1 NA 130 x 70 13.7 18.5 35.0 1
Oph-emb 15 I 231 330 0.17 0.78 8.1 144 4.2 NA 130 x 100 16.4 16.4 27.6 1
Oph-emb 24 Flat 251 560 0.29 0.84 6.3 144 4.5 NA 130 x 100 16.4 25.3 31.6 1
ODISEA C4 081 IIb 260 260 0.07 3.67 19.3 144 19.5 NA 280 x 190 33.2 14.2 22.1 7
Oph-emb 18 I 263 430 0.04 0.61 18.1 144 3.2 NA 130 x 100 16.4 5.5 19.2 1
ODISEA C4 033 Flat 271 540 0.204 12.8 25.4 144 67.8 NA 280 x 190 33.2 24.0 28.9 7
ISO-Oph-17 II 290c 1300 1.0 29.57 63.0 141 164.0 F 25 x 23 3.4 11.7 43.0 1, 8
Oph-emb 19 Flat 310 450 0.48 0.76 44.0 144 4.1 NA 130 x 100 16.4 3.2 35.8 1
Oph-emb 20 Flat 310 420 0.74 6.36 38.9 144 33.8 C 130 x 100 16.4 8.3 39.9 1
Oph-emb 17 Flat 328 390 1.61 44.25 47.8 132 280.0 F 34 x 25 3.8 18.3 48.4 1, 5, 9
ODISEA C4 107 Flat 332 332 0.03 2.6 29.8 144 13.8 NA 280 x 190 33.2 6.8 17.9 7
ISO-Oph 93 Flat 380 550 0.04 3.93 47.0 144 20.9 NA 130 x 100 16.4 5.4 19.2 1
Elias 2-27 Flat 410 1500 0.15 40.27 216.0 116 330.0 F 49 x 47 5.6 4.1 26.8 1, 7, 10, 11
Oph-emb 21d I 420 520 3.3 0.78 7.3 144 4.2 NA 130 x 100 16.4 19.0 58.0 1
Oph-emb 25 Flat 429 690 0.4 1.72 7.5 144 9.2 NA 130 x 100 16.4 34.3 34.2 1
CFHTWIR-Oph 43 Flat 440 570 0.12 0.38 4.4 144 2.0 NA 130 x 100 16.4 23.9 25.3 1
Oph-emb 23 Flat 440 570 0.12 8.68 50.8 144 46.2 NA 130 x 100 16.4 7.4 25.3 1
Oph-emb 26a Flat 500 620 0.97 1.9 27.8 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
Oph-emb 26b Flat 500 620 0.97 1.89 27.7 144 10.1 NA 130 x 100 16.4 6.3 42.7 1
F-MM7 Flat 570 690 0.72 1.17 27.0 144 6.2 NA 130 x 100 16.4 5.2 39.6 1
Oph-emb 27 Flat 570 690 0.72 13.0 24.9 144 69.0 NA 130 x 100 16.4 25.2 39.6 1
Oph-emb 28 Flat 610 720 1.1 2.13 10.3 144 11.3 NA 130 x 100 16.4 24.3 44.0 1
ODISEA C4 086 Flat 610 720 1.1 2.7 23.5 144 14.3 NA 280 x 190 33.2 9.1 44.0 7
ODISEA C4 045 Flat 613 613 0.35 2.87 54.7 144 15.2 NA 280 x 190 33.2 4.2 33.1 7
ODISEA C4 026 Flat 690 1000 0.0023 4.15 88.4 144 22.1 NA 280 x 190 33.2 3.5 9.4 7
ISO-Oph 37 Flat 710 810 0.24 20.92 91.0 141 116.0 NA 27 x 24 3.6 6.4 30.1 8
ODISEA C4 043 II 790 1600 0.092 1.74 32.8 144 9.2 NA 280 x 190 33.2 5.2 23.7 7
EM*SR 24S II 840 1300 1.3 21.83 58.0 115 182.0 C 37 x 33 4.0 10.7 45.9 8
ODISEA C4 144 II 980 1500 0.0011 1.48 57.1 144 7.8 NA 280 x 190 33.2 3.3 7.8 7
Elias 2-24 II 980 1500 1.9 59.05 115.0 136 352.0 F 37 x 34 4.8 8.6 50.5 1, 7, 10, 11
Elias 2-20 II 990 1600 0.5 17.96 54.0 138 104.0 F 32 x 23 3.7 10.4 36.2 1, 7, 10, 11
ODISEA C4 055 II 1000 1600 0.083 4.09 12.6 144 21.8 NA 280 x 190 33.2 29.7 23.1 7
ODISEA C4 070 II 1000 1600 0.22 7.66 36.0 144 40.7 NA 280 x 190 33.2 10.1 29.4 7
ODISEA C4 104 II 1000 1500 0.37 12.5 25.5 144 66.6 NA 280 x 190 33.2 23.5 33.5 7
WSB 52 II 1000 1500 0.37 11.24 27.0 136 67.0 F 33 x 27 4.1 19.7 33.5 7, 10, 11
ISO-Oph 2a II 1100 1600 0.11 11.85 72.0 144 63.0 C 30 x 21 3.6 6.1 24.8 1, 8
ISO-Oph 2b II 1100 1600 0.11 0.3 11.9 144 1.6 NA 130 x 100 16.4 5.9 24.8 1, 8
DoAr 20 II 1100 1400 0.89 7.3 41.5 144 38.8 F 130 x 100 16.4 8.3 41.8 1
EM* SR4 II 1100 1400 0.89 11.24 26.0 134 69.0 F 34 x 34 4.6 20.9 41.8 7, 10, 11
WSB 82 II 1100 1500 0.76 41.62 256.0 155 191.0 C 31 x 24 4.2 3.7 40.1 8
ISO-Oph 196 II 1200 1600 0.14 15.32 69.0 137 90.0 C 45 x 22 4.3 7.2 26.3 8
DoAr 44 II 1200 1500 0.75 11.6 60.0 146 60.0 C 34 x 25 4.3 7.2 40.0 8
ODISEA C4 060 II 1300 1700 0.13 2.9 30.3 144 15.4 NA 280 x 190 33.2 7.1 25.8 7
ODISEA C4 121 II 1300 1600 0.14 1.4 18.8 144 7.5 NA 280 x 190 33.2 8.0 26.3 7
ODISEA C4 118 II 1400 1700 0.17 2.55 21.0 144 13.6 NA 280 x 190 33.2 10.0 27.6 7
ODISEA C4 005 II 1400 1600 0.42 5.61 22.6 144 29.8 NA 280 x 190 33.2 15.4 34.6 7
ODISEA C4 018 II 1400 1600 0.13 2.33 14.5 144 12.4 NA 280 x 190 33.2 15.5 25.8 7
WaOph 6 II 1400 1700 1.2 22.09 87.0 123 161.0 F 58 x 54 6.9 6.8 45.0 1, 10, 11
RX J1633.9-2442 II 1500 1800 0.17 12.8 53.0 141 71.0 C 21 x 20 2.9 8.7 27.6 1, 8
DoAr 25 II 1500 1700 0.48 42.49 140.0 138 246.0 F 41 x 22 4.1 5.9 35.8 1, 7, 10, 11
DoAr 33 II 1600 1800 0.3 6.13 23.0 139 35.0 F 37 x 24 4.1 15.9 31.8 10, 11
ODISEA C4 017 II 1800 1900 0.085 0.98 28.6 144 5.2 NA 280 x 190 33.2 4.6 23.2 7
Chamaeleon I & II
TIC 454291385 I 74 74 1.0 0.65 22.2 189 2.0 NA 54 x 35 8.2 4.2 43.0 1, 5
IRAS 12500-7658 I 187 160 0.31 14.76 60.4 181 49.7 NA 149 x 81 19.9 8.1 32.1 1
V* GM Cha I 219 300 2.47 2.35 31.7 179 8.1 NA 120 x 90 18.6 6.1 53.9 1
IRAS 12553-7651 I 260 260 1.74 2.42 41.4 144 12.9 NA 130 x 100 16.4 6.0 49.4 1
IRAS 11030-7702 I 308 510 0.19 9.77 81.8 179 33.6 NA 74 x 43 10.1 5.0 28.4 1
ISO-ChaI 101 Flat 650 820 0.09 12.06 73.8 179 41.5 C 99 x 65 14.4 6.0 23.6 1
V* DK Cha II 660 1500 30.0 210.59 56.7 181 708.7 NA 146 x 82 19.8 68.2 100.6 1
V* HO Cha a Flat 710 900 0.97 4.35 20.4 179 15.0 NA 120 x 90 18.6 14.8 42.7 1
V* HO Cha b Flat 710 900 0.97 4.43 22.0 179 15.3 NA 120 x 90 18.6 13.5 42.7 1
ISO-ChaI 204 II 720 1100 0.18 0.46 11.9 179 1.6 NA 120 x 90 18.6 7.2 28.0 1
ISO-ChaI 207 II 1100 1500 0.09 8.75 90.7 179 30.1 NA 100 x 65 14.4 4.4 23.6 1
ISO-ChaI 237 II 1100 1600 0.37 8.9 25.8 179 30.6 NA 120 x 90 18.6 17.7 33.5 1
Ass Cha T 1-15 II 1400 1700 0.45 12.7 46.5 179 43.7 NA 100 x 65 14.4 10.1 35.2 1
Corona Australis
V* VV CrA A 0a 37 610 1.55 47.8 22.7 149 237.3 NA 110 x 80 14.0 94.5 48.0 1
IRS 5N 0 40 23 3.59 19.4 46.7 147 99.0 NA 50 x 30 5.7 10.7 59.2 1, 5
CrAus7-mm 0 54 310 37.75 1.54 16.2 149 7.7 NA 110 x 80 14.0 10.1 106.6 1
IRS 7A 0 54 310 37.75 3.05 8.7 149 15.2 NA 110 x 80 14.0 44.0 106.6 1
CrAus8-mm2 0 56 310 21.13 3.42 15.7 149 17.0 NA 110 x 80 14.0 18.2 92.2 1
IRS 7B-a 0 56 210 5.1 71.03 59.6 152 339.0 NA 54 x 42 7.2 18.7 64.6 1, 5
IRS 7B-b 0 56 210 5.1 7.54 24.1 152 36.0 NA 54 x 42 7.2 13.7 64.6 1, 5
IRAS 32 A 0 58 66 2.58 16.33 27.0 150 80.0 NA 35 x 23 4.3 20.4 54.5 1, 5
IRAS 32 B 0 58 66 2.58 8.98 22.9 150 44.0 NA 35 x 23 4.3 16.6 54.5 1, 5
SMM 2 I 72 15 1.09 45.38 126.5 149 225.4 C 110 x 80 14.0 6.8 43.9 1
V* V710 CrAe I 135 270 28.08 75.27 30.7 149 373.8 NA 120 x 80 14.6 82.5 99.0 1
IRS 5a Flat 208 130 2.71 0.99 4.5 149 4.9 NA 110 x 80 14.0 51.4 55.2 1
IRS 5b Flat 208 130 2.71 0.4 34.1 144 2.1 NA 130 x 100 16.4 3.1 55.2 1
IRS 2 I 235 390 12.99 97.57 74.3 149 484.6 C 110 x 80 14.0 22.0 81.6 1
V* S CrA A II 1000 1500 3.5 32.51 31.4 149 161.4 NA 110 x 80 14.0 36.9 58.8 1
V* S CrA B II 1000 1500 3.5 30.27 35.6 149 150.3 NA 110 x 80 14.0 28.2 58.8 1
Taurus
IRAS 04166+2706 0 61 61 0.4 15.67 18.3 156 71.0 NA 49 x 37 6.6 37.6 34.2 5
IRAS 04169+2702 I 163 163 1.5 22.29 28.6 156 101.0 C 48 x 37 6.6 23.8 47.6 5
TMC-1A I 183 183 2.3 31.32 30.1 137 184.0 NA 32 x 21 3.6 28.9 53.0 5
L1489 IRS I 213 213 3.4 17.59 485.2 146 91.0 C 105 x 78 13.2 2.0 58.4 5
HL Tau I 576 576 3.0 132.27 137.2 140 744.1 F 35 x 22 3.9 11.3 56.6 12, 13
BHR 71
BHR 71 IRS2 0 39 39 1.1 3.93 7.3 176 14.0 NA 70 x 50 10.0 56.0 44.0 5
BHR 71 IRS1 0 66 66 10.0 107.88 41.7 176 384.0 NA 72 x 53 10.9 48.1 76.5 5
Lupus I
IRAS 15398-3359 0 50 50 1.4 1.72 3.8 155 7.9 NA 43 x 36 6.1 87.1 46.8 5
101010 CFHTWIR-Oph 79 is also known as GY 263 or SKS 3-48 (Wilking et al. 2015; Narayanan et al. 2023). The classification of this source is debated. Narayanan et al. (2023) and Allen et al. (2002) classified this source as a Class II disk, while Wilking et al. (2015) classified it as a Flat spectrum disk based on J, H, and K band. The source is close to a Class I binary system, Oph-emb 14 VLA 1&2. In the CAMPOS survey, we assume the same bolometric temperature and class for all sources in the close multiple systems. a The SED and protostellar system properties are available in NASA/IPAC Infrared Science Archive. eHOPS classified the source as Class I, but given its bolometric temperature, it should be Class 0. b The eHOPS group calculated the spectral index between Spitzer/IRAC 4.5 $\mu$m and Spitzer/IRAC 24 $\mu$m, and follows Greene et al. (1994) for Class classification (Pokhrel et al. 2023). ODISEA C4 081 or eHOPS-oph-22 shows a negative slope between these wavelengths resembling a Class II by definition. However, the source also shows significant emission at longer wavelengths, resulting in a lower bolometric temperature. c In this paper, if the source has no corresponding eHOPS match, we adopted the non-extinction corrected bolometric temperature from Dunham et al. (2015). d eHOPS source labeled off center for IRAS $1-4\,\mu$m, eHOPS $T_{\rm bol}$ could be wrong, assigned values from Dunham et al. (2015). e eHOPS group updated the SED fit for this source recently, the original values published in Hsieh et al. (2024) for this source were outdated. eHOPS ${\rm T}_{\rm bol}$: Bolometric temperature from eHOPS catalog. The bolometric temperature is derived from SED fitting with Herschel Space Telescope longer wavelength data, but IR extinction correction is not applied. ${\rm L}_{\rm bol}$: Bolometric luminosity. Mass: Disk dust mass. R90%: Disk dust $R_{90\%}$ radius. dist: Distance of the source. $F_{1.3mm}$: 1.3 mm flux. Type: Type of the disk substructure. NA represents no disk substructure is detected. C represents a ring with a large central cavity. F represents the disk is centrally filled. Dun+2015 ${\rm T}_{\rm bol}$: Bolometric temperature from Dunham et al. (2015) catalog. The bolometric temperature is derived from SED fitting without Herschel Space Telescope longer wavelength data, but IR extinction correction is applied. $T_{\rm br,P}$: Brightness temperature of the disk calculated from the full Planck equation. $T_{\rm dust}$: Dust temperature of the disk adopting equation 1. Ref: References. References: (1) Sources observed in our CAMPOS survey. See also the CAMPOS data paper, Hsieh et al. (2024), for source identification and radius measurements. (2) Murillo et al. (2013) (3) Hsieh et al. (2020) (4) Maureira et al. (2022) (5) Ohashi et al. (2023) (6) Narayanan et al. (2023) (7) Cieza et al. (2019) (8) Cieza et al. (2021) (9) Segura-Cox et al. (2020) (10) Andrews et al. (2018a) (11) Huang et al. (2018) (12) ALMA Partnership et al. (2015) (13) Wu et al. (2018)
Table 2: Cross matching Corona Australis, Chamaeleon I & II, Ophiuchus North, and Ophiuchus sources in our CAMPOS survey
Name CAMPOS ID SSTgbs/SSTc2d Names eHOPS Name ISO Name IRAS sources
Ophiuchus & Ophiuchus North
IRAS 16442-0930 OphN-01-0 SSTgbs J1646582-093519 - - IRAS 16442-0930
IRAS 16459-1411 OphN-02-0 SSTgbs J1648456-141636 - - IRAS 16459-1411
/ WaOph 6
CB 68 SMM 1 OphN-03-0 SSTgbs J1657196-160923 - - IRAS 16544-1604
ISO-Oph 2a Oph-01-0 SSTc2d J162538.1-242236 - ISO-Oph 2 -
ISO-Oph 2b Oph-01-1 SSTc2d J162538.1-242236 - ISO-Oph 2 -
DoAr 20 Oph-02-0 SSTc2d J162556.1-242048 - ISO-Oph 6 IRAS 16229-2413
ISO-Oph 17 Oph-03-0 SSTc2d J162610.3-242055 - ISO-Oph 17 -
Elias 2-20 Oph-05-0 SSTc2d J162618.9-242820 - ISO-Oph 24 IRAS 16233-2421
Oph-emb 8 Oph-06-0 SSTc2d J162621.3-242304 eHOPS-oph-2 ISO-Oph 29 -
ISO-Oph 31 Oph-06-1 SSTc2d J162621.7-242250 - ISO-Oph 31 -
/ GSS 30 IRS3
DoAr 25 Oph-07-0 SSTc2d J162623.7-244314 - ISO-Oph 38 IRAS 16234-2436
Elias 2-24 Oph-08-0 SSTc2d J162624.1-241613 - ISO-Oph 40 IRAS 16233-2409
Oph-emb 9 Oph-09-0 SSTc2d J162625.5-242302 eHOPS-oph-4 - -
VLA 1623B Oph-10-0 SSTc2d J162626.4-242430 - - -
VLA 1623Ab Oph-10-1 SSTc2d J162626.4-242430 - - -
VLA 1623Aa Oph-10-2 SSTc2d J162626.4-242430 eHOPS-oph-5 - -
VLA 1623W Oph-10-3 SSTc2d J162625.6-242429 - - -
Oph-emb 22 Oph-11-0 SSTc2d J162640.5-242714 eHOPS-oph-7/8 ISO-Oph 54 -
IRAS 16237-2428 Oph-12-0 SSTc2d J162644.2-243448 eHOPS-oph-9 ISO-Oph 65 IRAS 16237-2428
Elias 2-27 Oph-13-0 SSTc2d J162645.0-242308 eHOPS-oph-10 ISO-Oph 67 -
Oph-emb 23 Oph-14-0 SSTc2d J162648.5-242839 - ISO-Oph 70 -
CFHTWIR-Oph 43 Oph-14-1 SSTc2d J162648.4-242835 - - -
DoAr 29 Oph-15-0 SSTc2d J162658.5-244537 - ISO-Oph 88 -
Oph-emb 21 Oph-16-0 SSTc2d J162702.3-243727 eHOPS-oph-12 ISO-Oph 92 -
ISO-Oph 93 Oph-17-0 SSTc2d J162703.0-242615 - ISO-Oph 93 -
Oph-emb 6 Oph-18-0 SSTc2d J162705.2-243629 eHOPS-oph-16 ISO-Oph 99 -
Oph-emb 20 Oph-19-0 SSTc2d J162706.8-243815 eHOPS-oph-17 ISO-Oph 103 -
Oph-emb 16 Oph-20-0 SSTc2d J162709.4-243719 eHOPS-oph-19 ISO-Oph 108 -
eHOPS-oph-20a Oph-21-0 - eHOPS-oph-20 ISO-Oph 121 -
eHOPS-oph-20b Oph-21-1 - eHOPS-oph-20 - -
eHOPS-oph-20c Oph-21-2 SSTc2d J162715.8-243843 eHOPS-oph-20 - -
Oph-emb 11 Oph-22-0 SSTc2d J162717.6-242856 eHOPS-oph-23 ISO-Oph 124 -
Oph-emb 28 Oph-23-0 SSTc2d J162721.5-244143 - ISO-Oph 132 -
Oph-emb 12 Oph-25-0 SSTc2d J162724.6-244103 eHOPS-oph-26 ISO-Oph 137 -
Oph-emb 14 VLA 1 Oph-26-0 SSTc2d J162726.9-244051 eHOPS-oph-29 ISO-Oph 141 IRAS 16244-2434
Oph-emb 14 VLA 2 Oph-26-1 SSTc2d J162726.9-244051 - ISO-Oph 141 IRAS 16244-2434
CFHTWIR-Oph 79 Oph-26-2 - - - IRAS 16244-2432
Oph-emb-13 Oph-27-0 SSTc2d J162728.0-243933 eHOPS-oph-30 ISO-Oph 143 -
Oph-emb 19 Oph-28-0 SSTc2d J162728.4-242721 - ISO-Oph 144 -
Oph-emb 26a Oph-29-0 SSTc2d J162730.2-242743 - ISO-Oph 147 -
Oph-emb 26b Oph-29-1 SSTc2d J162730.2-242743 - ISO-Oph 147 -
Oph-emb 24 Oph-30-0 SSTc2d J162737.2-244238 eHOPS-oph-35 ISO-Oph 161 -
Oph-emb 27 Oph-31-0 SSTc2d J162739.8-244315 - ISO-Oph 167 IRAS 16246-2436
F-MM7 Oph-31-1 SSTc2d J162739.8-244315 - ISO-Oph 167 IRAS 16246-2436
Oph-emb 1 Oph-32-0 SSTc2d J162821.6-243623 eHOPS-oph-42 - IRAS 16253-2429
Oph-emb 18 Oph-33-0 SSTc2d J162857.9-244055 eHOPS-oph-43 - -
Oph-emb 17 / IRS 63 Oph-34-0 SSTc2d J163135.6-240129 eHOPS-oph-44 - IRAS 16285-2355
Oph-emb 4 Oph-35-0 SSTc2d J163136.8-240420 eHOPS-oph-46 - -
Oph-emb 25 Oph-36-0 SSTc2d J163143.8-245525 eHOPS-oph-47 ISO-Oph 200 -
Oph-emb 15 Oph-38-0 SSTc2d J163152.5-245536 eHOPS-oph-48 ISO-Oph 203 -
Oph-emb 10a Oph-39-0 SSTc2d J163201.0-245643 eHOPS-oph-49 ISO-Oph 209 -
Oph-emb 10b Oph-39-1 SSTc2d J163201.0-245643 - ISO-Oph 209 -
IRAS 16293-2422A Oph-40-0 - eHOPS-oph-51 - IRAS 16293-2422
IRAS 16293-2422B Oph-40-1 SSTc2d J163222.6-242832 - - -
EDJ 1013 Oph-41-0 SSTc2d J163355.6-244205 - - -
Chamaeleon I & II
IRAS 11030-7702 ChamI-01-0 SSTgbs J1104227-771808 eHOPS-cha-2 ISO-ChaI 46 IRAS 11030-7702
TIC 454291385 ChamI-02-0 SSTgbs J1106464-772232 eHOPS-cha-3 - -
/ Ced 110 IRS4B
ISO-ChaI 101 ChamI-05-0 SSTgbs J1107213-772211 - ISO-ChaI 101 -
Ass Cha T 1-15 ChamI-06-0 SSTgbs J1107435-773941 - ISO-ChaI 112 -
V* HO Cha a ChamI-07-0 - - ISO-ChaI 126 -
V* HO Cha b ChamI-07-1 SSTgbs J1108029-773842 - ISO-ChaI 126 -
V* GM Cha ChamI-08-0 SSTgbs J1109285-763328 eHOPS-cha-7 ISO-ChaI 192 -
ChamI-9 mm ChamI-09-0 - - - -
ISO-ChaI 204 ChamI-09-1 SSTgbs J1109461-763446 - ISO-ChaI 204 -
ISO-ChaI 207 ChamI-10-0 SSTgbs J1109472-772629 - ISO-ChaI 207 -
IR Cha INa4 ChamI-11-0 SSTgbs J1110033-763311 eHOPS-cha-11 - -
ISO-ChaI 237 ChamI-12-0 SSTgbs J1110113-763529 - ISO-ChaI 237 -
TIC 454329229 ChamI-13-0 SSTgbs J1111107-764157 - - -
V* DK Cha ChamII-01-0 SSTc2d J125317.2-770710 - - IRAS 12496-7650
IRAS 12500-7658 ChamII-02-0 SSTc2d J125342.9-771511 eHOPS-cha-12 - IRAS 12500-7658
IRAS 12553-7651 ChamII-03-0 STc2d J125906.6-770740 eHOPS-cha-13 ISO-ChaII 28 IRAS 12553-7651
Corona Australis
V* S CrA B CrAus-01-0 - - ISO-CrA 116 IRAS 18577-3701
V* S CrA A CrAus-01-1 SSTgbs J1901086-365720 - ISO-CrA 116 IRAS 18577-3701
IRS 2 CrAus-02-0 SSTgbs J1901415-365831 eHOPS-cra-2 - -
IRS 5a CrAus-03-0 SSTgbs J1901480-365722 - - -
IRS 5b CrAus-03-1 SSTgbs J1901480-365722 eHOPS-cra-3 - -
IRS 5N CrAus-04-0 SSTgbs J1901484-365714 eHOPS-cra-4 - -
V* V710 CrA CrAus-05-0 SSTgbs J1901506-365809 eHOPS-cra-6 - -
IRS 7A CrAus-07-0 SSTgbs J1901553-365721 eHOPS-cra-7 - -
SMM1C CrAus-07-1 - - - -
CrAus7-mm CrAus-07-2 - - - -
CrAus8-mm1 CrAus-08-0 - - - -
IRS 7B-a CrAus-08-1 SSTgbs J1901564-365728 eHOPS-cra-8 - -
CXO 34 CrAus-08-2 J190155.$7$
IRS 7B-b CrAus-08-3 SSTgbs J1901564-365728 - - -
SMM 2 CrAus-09-0 SSTgbs J1901585-365708 eHOPS-cra-9 - -
IRAS 32 A CrAus-10-0 SSTgbs J1902586-370735 eHOPS-cra-10 ISO-CrA 182 IRAS 18595-3712
IRAS 32 B CrAus-10-1 - - ISO-CrA 182 IRAS 18595-3712
V* VV CrA A CrAus-11-0 SSTgbs J1903068-371249 eHOPS-cra-11 - IRAS 18597-3717
V* VV CrA B CrAus-11-1 - - - IRAS 18597-3717
 [^11] CAMPOS ID: Sources observed in our CAMPOS survey. See also the CAMPOS data paper (Hsieh et al. 2024). SSTgbs/SSTc2d Names: Spitzer Space Telescope Gould Belt Survey / Space Telescope core to disk Survey Names. See (Dunham et al. 2015) for the spectral energy distribution (SED) fitting and the bolometric temperature, and the bolometric luminosity of the source. eHOPS Name: Extension of HOPS Out to 500 ParSecs (eHOPS) catalog, represents the latest and most reliable SED fitting to date. The eHOPS catalog contains 1-850 $\mu$m SEDs assembled from 2MASS, Spitzer, Herschel, WISE, and JCMT/SCUBA-2 data. The first paper of eHOPS for Serpens and Aquila molecular clouds was published by Pokhrel et al. (2023). For all other clouds, the SED and protostellar system properties are available in NASA/IPAC Infrared Science Archive. ISO Name: Infrared Space Observatory (ISO) source name. IRAS Name: Infrared Astronomical Satellite (IRAS) source name.
Table 3: Cross Matching ODISEA Survey with eHOPS and SSTgbs/SSTc2d
ODESIA ID SSTgbs/SSTc2d Names eHOPS Name Other Name Included? Notes
ODISEA C4 005 J162218.5-232148 - - Y -
ODISEA C4 017 J162506.9-235050 - - Y -
ODISEA C4 018 J162524.3-242944 - - Y -
ODISEA C4 022 J162538.1-242236 - ISO-Oph-2 Y -
ODISEA C4 026 J162546.6-242336 - - Y -
ODISEA C4 027 J162556.1-242048 - EM* SR4 Y -
ODISEA C4 030 J162610.3-242055 - ISO-Oph-17 Y -
ODISEA C4 033 J162617.2-242345 eHOPS-oph-1 - Y -
ODISEA C4 034 J162618.9-242820 - Elias 2-20 Y -
ODISEA C4 038 J162623.5-242439 - ISO-Oph 37 Y -
ODISEA C4 039 J162623.6-244314 - DoAr 25 Y -
ODISEA C4 041 J162624.0-241613 - Elias 2-24 Y -
ODISEA C4 042 J162625.4-242301 eHOPS-oph-4 Oph-emb 9 Y -
ODISEA C4 043 J162627.5-244153 - - Y -
ODISEA C4 045 - eHOPS-oph-6 - Y -
ODISEA C4 047 J162640.4-242714 eHOPS-oph-7 ISO-Oph-54 Y -
ODISEA C4 051 J162645.0-242308 eHOPS-oph-10 Elias 2-27 Y -
ODISEA C4 055 J162648.9-243825 - - Y -
ODISEA C4 057 - eHOPS-oph-11 - N unresolved
ODISEA C4 060 J162656.7-241351 - - Y -
ODISEA C4 062 J162658.5-244537 - EM*SR 24S Y -
ODISEA C4 064 J162703.5-242005 eHOPS-oph-13/14 - Y -
ODISEA C4 065 - eHOPS-oph-15 - N unresolved
ODISEA C4 067 - eHOPS-oph-16 Oph-emb 6 Y -
ODISEA C4 068 - eHOPS-oph-17 Oph-emb 20 Y -
ODISEA C4 070 J162709.0-243408 - - Y -
ODISEA C4 071 - eHOPS-oph-18 - N unresolved
ODISEA C4 080 - eHOPS-oph-21 - N unresolved
ODISEA C4 081 - eHOPS-oph-22 - Y -
ODISEA C4 082 - eHOPS-oph-23 Oph-emb 11 Y -
ODISEA C4 086 J162721.4-244143 - - Y -
ODISEA C4 087 - eHOPS-oph-24 - N unresolved
ODISEA C4 090 - eHOPS-oph-26 Oph-emb 12 Y -
ODISEA C4 091 - eHOPS-oph-28 - N unresolved
ODISEA C4 099 - eHOPS-oph-35/36 Oph-emb 24 Y -
ODISEA C4 102 - eHOPS-oph-37 - Y -
ODISEA C4 104 J162739.4-243915 - - Y -
ODISEA C4 105 J162739.8-244315 - - N unresolved
ODISEA C4 107 - eHOPS-oph-38 - Y -
ODISEA C4 111 - eHOPS-oph-40 - N unresolved
ODISEA C4 112 - eHOPS-oph-41 - N unresolved
ODISEA C4 114 J162816.5-243658 - ISO-Oph 196 Y -
ODISEA C4 118 J162854.0-244744 - - Y -
ODISEA C4 119 - eHOPS-oph-43 Oph-emb 18 Y -
ODISEA C4 121 J163023.3-245416 - - Y -
ODISEA C4 127 J163133.4-242737 - DoAr 44 Y -
ODISEA C4 130 J163135.6-240129 eHOPS-oph-44 Oph-emb 17 Y -
ODISEA C4 131 - eHOPS-oph-46 Oph-emb 4 Y -
ODISEA C4 132 J163143.7-245524 eHOPS-oph-47 Oph-emb 25 Y -
ODISEA C4 135 - eHOPS-oph-48 Oph-emb 15 Y -
ODISEA C4 141 J163355.6-244205 - RX J1633.9-2442 Y -
ODISEA C4 143 J163945.4-240203 - WSB 82 Y -
ODISEA C4 144 J163952.9-241931 - - Y -
 [^12] SSTgbs/SSTc2d Names: Spitzer Space Telescope Gould Belt Survey / Space Telescope core to disk Survey Names. See (Dunham et al. 2015) for the spectral energy distribution (SED) fitting and the bolometric temperature, and the bolometric luminosity of the source. eHOPS Name: Extension of HOPS Out to 500 ParSecs (eHOPS), is the latest, most reliable SED fitting to date. The eHOPS catalog contains 1-850 $\mu$m SEDs assembled from 2MASS, Spitzer, Herschel, WISE, and JCMT/SCUBA-2 data. The first paper of eHOPS for Serpens and Aquila molecular clouds was published by Pokhrel et al. (2023). For all other clouds, the SED and protostellar system properties are available in NASA/IPAC Infrared Science Archive. Other Name: Names shown in Table 1, other higher resolution data available.
A.1 Bolometric temperature as a proxy for relative evolutionary age
Tracing protostellar disk age is extremely difficult. For embedded protostars, the luminosity is dominated by accretion, resulting in a large spread in observed luminosity. This means individual protostellar ages cannot be directly determined from traditional isochrome fitting. Instead, the age sequence for protostellar disks is based on the dissipation of the protostellar envelope, as indicated by observables such as bolometric temperature ($T_{\rm bol}$) (Chen et al. 1995; Myers & Ladd 1993).
$T_{\rm bol}$ is the temperature of a blackbody having the same flux-weighted mean frequency as the observed continuum spectrum (Chen et al. 1995). As a protostar evolves toward the main sequence, the circumstellar envelope dissipates, shifting the $T_{\rm bol}$ to a higher temperature. We followed the standard method (Wilking 1989; Dunham et al. 2015) to estimate the age for the onset of the protostellar disk substructures (at $T_{\rm bol}=200$, $400$ K). The standard approach for determining the duration of each evolutionary class involves calculating the ratio of the number of sources in the evolutionary stage of interest to the number in a reference stage, and then multiplying this ratio by the duration of the reference evolutionary class (Dunham et al. 2015; Wilking et al. 1989; Evans et al. 2009; Dunham et al. 2014). Following Evans et al. (2009), Dunham et al. (2014) and Dunham et al. (2015), we adopted Class II as the reference evolutionary class and assumed a Class II duration of 2 million years. We used the Spitzer Space Telescope cores to disks (c2d) and Gould Belt Legacy survey (gbs) catalog of 3000 young stellar objects in the 18 molecular clouds to estimate the age (Dunham et al. 2015). We divided the number of protostars with bolometric temperatures less than $200$ and 400400400\,400K by the total number of protostars in Class 0, Class I, flat-spectrum phase, and Class II phase combined. We multiplied this ratio by 2 million years to obtain the corresponding age of 0.2–0.3 Myr, and 0.3–0.4 Myr, at $T_{\rm bol}=200,400$ K, respectively.
Note that this commonly used number-counting method does not take into account different star-formation rates in different molecular clouds, assumes a steady state constant star formation rate, and treats the duration of each stage as a single value rather than a distribution (Kristensen & Dunham 2018). Evans et al. (2009) developed the idea that the protostellar lifetimes should be thought of as half-lives, in which they represent the timescales at which half of the objects will have moved through and left that $T_{\rm bol}$ range. Kristensen & Dunham (2018) accounting for non-steady state star formation rates and adopted a sequential nuclear decay model to estimate the duration of each protostellar stage, and found that the half-life of Class 0 and Class I are $\sim 4.7\times 10^{4}$ and $\sim 8.7\times 10^{4}$ years, respectively.
If we adopt the duration from Kristensen & Dunham (2018), the onset of disk substructure will occur much earlier. In this paper, we adopt a more conservative number-counting method for the age estimate. We also emphasize that the age estimate for the onset of protostellar disk substructures is only a zeroth-order estimate, and future work is needed to improve the $T_{\rm bol}$-age relation.